Tohban Report 2009-12-02
From RHESSI Wiki
|Start Date:||25 November 2009|
|End Date:||02 December 2009|
|Next Tohban:||Pascal Saint-Hilaire|
|List all reports|
Two small active regions (11032 and 11033) rotated off the disk over the weekend, but neither produced any significant activity; no RHESSI or GOES flares were recorded. Activity is minimal at the moment.
The SSR has been trending unusually high, getting up to ~70% early this week. This is most likely due to low-level noise from G2; no fix has yet been implemented, but David Smith is formulating a solution. As a stopgap, front segment data during eclipse was disabled starting Tuesday, to be re-enabled once the fill level zeros out.
DATA GAPS FOR 2009/11/26 GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 GAP START TIME GAP END TIME GAP (SEC) 2009-11-26T06:00:00.000 -- 2009-11-26T06:15:00.000 900.00000
A spin-up was performed on Tuesday, with no anomalies.